Keeley:

A website-hosting service a one-time setup fee of $350 plus d dollars for each month. If a business owner paid $1010 for the first 12 months, including the setup fee, what is the value of d?

Keeley:

I need help with this one.

Imagine:

350+d=1010

Keeley:

Thanks!

Imagine:

1010-350=660/7

Imagine:

660/12*=55=d

Imagine:

55=d

Imagine:

All done!

Keeley:

Thanks!

Imagine:

You're welcome!